OpenStudy (anonymous):

Find two numbers such that their sum is 12 and their product is 29

OpenStudy (gorv):

x+y=12 xy=29 y=29/x x+29/x=12 (x^2+29)/x=12 x^2+29=12x x^2-12x+29=0

OpenStudy (gorv):

now factor it

OpenStudy (gorv):

@rhagemore